Solve the equation for x.(2x – 1)2 – 9 = 0
Answer by sofiyac(983)

you need to make sure you copy the problem correctly
(2x – 1)^2 – 9 = 0 first add 9 to each side of the equal sign
(2x – 1)^2 = 9 next take square root of each side, you get two cases due to the square root since -3*-3=9 AND 3*3=9
case 1:
2x-1=3 add 1 to each side of the equal sign
2x=4 divide each side by 2
x=2
case 2:
2x-1=-3 add 1 to each side of the equal sign
2x=-2 divide each side by 2
x=-1
so the solutions are x=2 or x=-1